OpenStudy (anonymous):

how do you write X Cannot Equal 6 in interval notation?

jimthompson5910 (jim_thompson5910):

Take the interval \[\Large (-\infty, \infty) \] which is the set of all real numbers and "poke a hole in it" so to speak to get this \[\Large (-\infty, 6) \cup (6, \infty) \]

jimthompson5910 (jim_thompson5910):

another way to see it: you're taking the set of numbers that's less than 6 and combining it with the set of numbers that are greater than 6 we're not including the actual value of 6

OpenStudy (anonymous):

Okay thank you so much (:

jimthompson5910 (jim_thompson5910):

you're welcome